The price of pens today is N10, and the quantity demanded is 1m. Next year the price rises to N11 and the quantity demanded falls to 950,000. What is the price elasticity of demand and what type of elasticity is this? *


1
Expert's answer
2021-04-06T07:06:13-0400

We can find the price elasticity of demand as follows:


"E_d=\dfrac{\%\Delta Q}{\%\Delta P}=\dfrac{\dfrac{Q_2-Q_1}{0.5(Q_2+Q_1)}}{\dfrac{P_2-P_1}{0.5(P_2+P_1)}},"

"\u200b\t\n \n\u200b\t\n \n\u200b"

"E_d=\dfrac{\dfrac{950000-1000000}{0.5(950000+1000000)}}{\dfrac{N11-N10}{0.5(N11+N10)}}=-0.54"


As we can see, the price elasticity of demand equals 0.54. Since "E_d<1" the demand is inelastic.